OpenStudy (anonymous):

Using complete sentences, explain how to convert your birthday function into standard form.

OpenStudy (anonymous):

My birthday function: f(x) = 2(x-11)^2 + 27

OpenStudy (phi):

standard form is y(x) = a x^2 + b x + c where a, b, and c are numbers. the first step is "expand" (x-11)^2 = (x-11)(x-11) (you should have born in january.. you would have an easier problem)

OpenStudy (anonymous):

lol sorry for not being born in january xD

OpenStudy (anonymous):

So first we expand it to (x-11)^2 = (x-11)(x-11)?

OpenStudy (phi):

use FOIL (for example) to multiply it out. combine like terms. add in the +27 and you will have standard form

OpenStudy (anonymous):

so its like -11*-11 = 121 like four times lol?

OpenStudy (anonymous):

I'm confused because theres like 3 terms instead of four? or does the exponent count as one?

OpenStudy (phi):

The first step is to multiply (x-11)(x-11) the video will explain it faster (about 3 minutes) than we can do it by typing here...

OpenStudy (anonymous):

@phi Can you help me? I need help on a question

OpenStudy (anonymous):

(x-11)^2 = (x-11)(x-11) so (x-11)(x-11) x * x + x * (-11) + (-11) * x + (-11) * (-11)?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes. x*x is written in short-hand as x^2 (x^2 means x*x) x^2 + x * (-11) + (-11) * x + (-11) * (-11) x*(-11) is written -11x (order does not matter when multiplying and (-11)*x is written -11x (algebra leaves out the * sign) of course -11*-11= 121 x^2 - 11x - 11 x + 121 now combine the "like terms" -11 x's and take away another 11 x's , how many x's ?

OpenStudy (anonymous):

so after FOIL my equation should of been x^2 - 11x - 11 x + 121?

OpenStudy (phi):

you still should combine -11x - 11x

OpenStudy (anonymous):

x^2 - 11x - 11 x + 121 combine like terms... x^2 - 11x^2 + 121?

OpenStudy (phi):

you only get x^2 by multiplying maybe its clearer if we had x + x if you have 1 x and add another x you get 2 x's or 2x it works the same with negative numbers x - 2x if you have 1 x and take away 2 x's you get -1 x -11x - 11 x is the same idea: -11 x's take away 11 x's = -22 x

OpenStudy (anonymous):

Alright I gotcha

OpenStudy (anonymous):

so x^2 - 22x + 121?

OpenStudy (phi):

yes. we started with f(x) = 2(x-11)^2 + 27 you expanded (x-11)^2 = (x-11)(x-11) = x^2 - 22x + 121 replace (x-11)^2 in your formula with the expanded version (be sure to use parens) f(x)= 2(x^2 - 22x + 121) + 27 next step is distribute the 2. that means multiply every term inside the parens by 2

OpenStudy (anonymous):

f(x)= 2(x^2 - 22x + 121) + 27 f(x) =(2x^2 - 44 + 242 + 54)?

OpenStudy (phi):

first the x does not disappear in 2*(-22)x second, the 27 is not part of (x-11)^2. You leave it alone.

OpenStudy (phi):

remember we are doing 2(x-11)^2 which is the same as 2(x^2 - 22x + 121) the +27 is added in later

OpenStudy (anonymous):

my bad on the 44 without x, I didnt see the x earlier f(x)= (2x^2 - 44x + 242) + 27

OpenStudy (anonymous):

@Brad1996 Check out this thread: http://openstudy.com/study#/updates/52570bd1e4b0ede9e447371c

OpenStudy (phi):

last step, we can drop the parens. (there is nothing "out front" multiplying them, so it is ok) f(x)= 2x^2 - 44x + 242+ 27 last step simplify by "combining like terms"... in this case the last two numbers can be added together.

OpenStudy (phi):

yes, looks good